what is the slope of the line with the equation x+2y=5
Answer by Alan3354(69443)

what is the slope of the line with the equation x+2y=5
--------------
Slope, m = -1/2
To find the slope, put the eqn in the slope-intercept form, y = xm + b
All that means is solve for y.
x+2y=5
2y = -x + 5
y = (-1/2)x + 5/2
m = -1/2